# Benchmarking MedMNIST dataset on real quantum hardware

### v1

@@ 2.4 Inference on real IBM Cleveland hardware
The trained quantum models demonstrating the best classification performance on the simulator were selected for validation on the IBM Cleveland quantum hardware. During the inference stage of our QML framework, the trained quantum models with optimized parameters were utilized for evaluating the test samples. For each testing sample, the corresponding quantum circuit is dynamically constructed and transpiled at optimization level 3 to minimize circuit depth and gate count while ensuring compatibility with the hardware topology. Moreover, qubit connectivity was carefully configured to align with the hardware’s topology, minimizing the need for SWAP gates, which can introduce unnecessary noise and overhead. To further enhance the quantum model performance, we employed error suppression techniques, i.e. dynamical decoupling and gate twirling (Twir) during the sampling process on real IBM hardware. The DD is used to suppress the decoherence effects introduced due to the unwanted interaction between the qubits and the environment. Gate twirling applies random single-qubit gates around a target gate to average out coherent errors. The sampling on real hardware is performed using Qiskit SamplerV2 in Session execution mode. Moreover, we also integrated M3 error mitigation [42] in our QML workflow. M3 mitigation is a scalable quantum error correction approach that operates within the noisy bitstring subspace, correcting correlated and uncorrelated errors using a memory-efficient, matrix-free iterative solver. The integration of these error suppression and mitigation strategies enables quantum models to achieve optimal inference performance on real quantum hardware. Lastly, classical post-processing is applied on the mitigated bitstrings to obtain the predicted classification labels. To achieve this, the mitigated bitstrings are processed to compute qubit probabilities defined as:
@@
where $\langle Z_{q}\rangle$ denotes the expectation value of the Pauli-Z operator for the $q$-th qubit, calculated as the difference between the probabilities of measuring 0 and 1. This expectation value vector is obtained for each test sample is processed to obtain the predicted class labels. The results from all test samples were aggregated to compute the performance metrices, providing insights into the performance of the quantum circuits on real quantum hardware. By conducting inference on real quantum device, we demonstrated the feasibility of deploying noise-resilient and device-aware quantum circuits for practical medical image classification tasks.
@@ 3 Experimental results
In this section, we conducted extensive experimental analysis to evaluate our QML workflow using both simulator and IBM Cleveland real hardware to benchmark MedMNIST datasets [1]. The MedMNIST contains 12 two-dimensional biomedical imaging datasets of different medical modalities and complexity levels. These 12 datasets includes 2-class as well as multi-class datasets including 4-class, 5-class, 7-class, 8-class, 9-class, and 11-class. For this benchmarking, we considered 8 datasets out of 12 by considering atleast one dataset from each category of different classes as shown in Table 1. The different medical modalities considered in these 8 datasets are illustrated in Fig. 2. We considered two widely used classification metrices i.e., accuracy (ACC) and area under the curve (AUC) to measure the performance of QML models. It is important to note that ACC provides a straightforward interpretation of model performance but it is highly sensitive to class imbalance, often leading to biased evaluations. On the contrary, AUC offers a more comprehensive evaluation of the classifier’s separability and is inherently robust to class imbalance. This makes AUC particularly suitable for datasets where class distributions are skewed. It is clear from the Table 1 that MedMNIST datasets are quite imbalance in nature. Therefore, we prioritized AUC over ACC to select model for real hardware execution. By focusing on AUC, we ensured that the chosen quantum models were able to effectively distinguish between classes across various thresholds, providing a fairer and more reliable assessment of performance.
@@
The classical post-processing of M3-mitigated bitstrings in the inference stage provides expectation values vector for all test samples. These expectation values are used to calcuate the ACC and AUC metrices. For binary datasets (where we measure one qubit), the accuracy is calculated by measuring the proportion of predictions that deviate from the true labels by less than a threshold of 1. It evaluates how many predictions are close enough to the actual labels, providing an overall measure of correctness for the model’s outputs in binary classification tasks. The binary AUC is computed by assessing the separability of the two classes based on the predicted expectation values. We generates signed prediction labels to align with the true labels and calculates the AUC score using the AUC score function [43] from scikit-learn. For the multi-class datasets (where we are measuring more than 2 qubits), the model’s predicted class is determined by taking the index of the maximum value along each row of expectation values vector. This is compared against the true labels to compute the classification accuracy as:
@@
To calculate the AUC metric for multi-class datasets, we first applied softmax activation function to convert the raw expectation values into class probabilities. This ensures that the outputs for each test sample sum to 1 and can be interpreted as probabilities for each class. Then, we use the AUC function [43] from scikit-learn to calculate the AUC.
Table 2: Performance analysis of QML framework on simulator based on classification accuracy (ACC) and area under the curve (AUC) metrices under different configurations of quantum models.
@@
Table 3: Performance analysis of QML framework on simulator based on classification accuracy (ACC) and area under the curve (AUC) metrices under different configurations of quantum models.
@@
We present the noiseless results of the QML framework using simulator based on ACC and AUC metrices under different settings of input embedding features and quantum circuit trainable parameters for the considered MedMNIST datasets in Tables 2 and 3. The QML evaluation is performed by considering all the training and testing samples of MedMNIST datasets as given in Table 1. The bold values in the Tables 2 and 3 represent the best accuracy and AUC values. For PneumoniaMNIST dataset, our QML model with 4 qubits, 49 input embedding features, and 120 trainable parameters provides the highest classification accuracy of 85.26% and AUC of 81.97% as compared to the other settings as shown in Table 2. In the case of BreastMNIST dataset, we got best accuracy of 81.40% for the 4-qubit quantum model having 49 embedding features and 100 trainable parameters, while highest AUC of 72.20% is achieved for the model with 64 embedding features and 80 trainable parameters. Similarly, we got the highest accuracy of 44.30% and AUC of 65.90% for 8-qubit quantum model having 49 embedding features and 100 trainable parameters for OCTMNIST dataset. For RetinaMNIST dataset, we achieved best classification accuracy of 51.75% for the 5-qubit quantum model having 64 input embedding features and 120 trainable parameters. The highest AUC of 66.48% is obtained for the quantum model having 64 embedding features and 60 trainable parameters as depicted in Table 2. A 7-qubit quantum model with 64 input embedding features and 120 trainable parameters provides highest accuracy of 67.38% and AUC of 73.43% for the 7-class DermaMNIST dataset as shown in Table 3. For a 8-class BloodMNIST dataset, we achieved the highest accuracy of 48.20% and AUC of 82.10% using the QML model with 64 embedding features and 120 trainable parameters. Similarly, we got the best accuracy of 37.23% and AUC of 78.40% for a 9-class PathMNIST dataset using the QML model having 64 embedding features and 120 trainable parameters. Lastly, the QML model with 64 features and 100 trainable parameters provides best classification accuracy of 37.86% for 11-class OrganSMNIST dataset, while best AUC of 78.02% is achieved for the QML model having 64 embedding features and 120 trainable parameters. It is clear from these results that the quantum models without any classical neural networks part and with very limited input features showing promising capability to classify medical imaging datasets. Also, it is observed from the Tables 2 and 3 that there is a significant difference between the accuracy and AUC values, for instance, we got an accuracy of 48.20% and AUC of 82.10% for BloodMNIST dataset. Similar trend can be seen for the PathMNIST and OrganSMNIST datasets. Note that we considered AUC as our primary performance evaluation metric because it is less sensitive to the class imbalance of MedMNIST datasets as compared to the accuracy metric. Moreover, it can be noted that for most of the multi-class datasets, we got best results corresponding to the QML models having more number of embedding features and trainable parameters in the quantum circuit.
@@
We also conducted an ablation study on IBM Cleveland real hardware to understand the impact of error suppression and error mitigation techniques on the performance of QML models for medical imaging classification as shown in Table 4. We performed this ablation on the OCTMNIST dataset because of the significant effect of hardware noise on the performance of QML model. For OCTMNIST dataset, our QML workflow provides the noiseless accuracy of 44.30% and AUC of 65.90% as shown in Table 2. When we run this quantum model on real IBM hardware without any error suppression and mitigation techniques, we got an accuracy of 28.60% and AUC of 54.50%. Therefore, we observed the reduction of 15.70% and 11.40% in accuracy and AUC, respectively due to the noise of IBM Cleveland hardware. In this ablation, since we are using sampler in the inference stage as shown in Fig. 1, we considered two error suppression techniques i.e., dynamical decoupling and gate twirling and one error mitigation technique i.e., M3 mitigation. It is observed from the Table 4 that when we use both error suppression techniques and M3 mitigation together (DD+Twirling+M3), we got best accuracy of 39.70% and AUC of 62.20%, thereby providing an improvements of 11.10% and 7.7% in accuracy and AUC respectively as compared to the case when no suppression and mitigation is used. Based on this ablation study on OCTMNIST dataset, we used this combination (DD+Twirling+M3) for all the considered MedMNIST datasets.
@@
We selected the best performing QML models from Tables 2 and 3 based on the AUC metric for inference on real IBM Cleveland hardware for all the considered MedMNIST datasets. Due to the hardware limitations, we considered 1000 testing samples to evaluate the QML models for the datasets having large number of testing samples. We presented noiseless and IBM hardware results (with and without error suppression and mitigation) together in Table 5 for better understanding along with the quantum models specifications such as number of qubits, number of embedded features/gates, trainable parameters, quantum circut gates, and circuit depth. It is observed from the Table 5 that even with the small number of input embedding features, our QML models consistently provide promising results in terms of classification accuracy and AUC metrices. It is worth noting that the noiseless results and hardware results (without error suppression and mitigation) are quite comparable for most of the MedMNIST datasets except OCTMNIST and OrganSMNIST dataset. This is because we generated device-aware quantum circuits with optimal circuit gates, depth, and robust against hardware noises. For OCTMNIST dataset, the classification accuracy and AUC on the hardware (without error suppression and mitigation) reduce by 15.70% and 11.40% respectively as compared to the noiseless results. Similarly, accuracy and AUC decrease by 14.11% and 4.8% respectively for OrganSMNIST dataset as compared to the noiseless results. This indicates that there is scope for further improvement in the design of device-aware quantum circuits. To mitigate the effect of hardware noises, we run the hardware experiments for all the considered MedMNIST datasets using error suppression and error mitigation combination (DD+Twirling+M3) based on the ablation study conducted in Table 4. We noted improvements in classification accuracy and AUC results using DD+Twirling+M3 combination as compared to the hardware results without error suppression and mitigation for all the datasets except BreastMNIST dataset. For instance, the classification accuracy and AUC values are significantly improved by 11.10% and 7.7% respectively as compared to the hardware results without error suppression and mitigation for OCTMNIST dataset. For BreastMNIST dataset, the accuracy and AUC values decreased by 1.3% and 2.4% respectively as compared to the hardware results without error suppression and mitigation.
@@
It is important to mention that all classical ML models leverage the full spatial dimensions of input images, whereas our QML model operates on a significantly reduced feature set, dictated by quantum hardware constraints. Despite this, our QML model achieves promising results on real IBM cleveland hardware in terms of classification accuracy and AUC values. For the 2-class PneumoniaMNIST dataset, we got very good classification accuracy and AUC of 85.42% and 82.18%, respectively. Similarly, we achieved an accuracy and AUC of 76.90% and 71.40%, respectively in the case of 2-class BreastMNIST dataset. Even for 7-class DermaMNIST dataset, our QML model obtained very good results with an accuracy and AUC values of 67.70% and 69.82%, respectively. Our QML model is able to achieve the AUC of more than 70% for all considered MedMNIST datasets except OCTMNIST, RetinaMNIST, and DermaMNIST datasets. It is worth mentioning that for 2-class PneumoniaMNIST and 8-class BloodMNIST datasets, we got the AUC of more than 80% on real quantum hardware. To the best of our knowledge, this benchmarking study is the first to provide a comprehensive evaluation of QML (without using any classical neural networks) on real IBM quantum hardware considering large-scale MedMNIST datasets. Our study provides a critical evaluation of QML for realistic applications, showcasing its capability even in the evolving phase of quantum computing. By systematically benchmarking QML performance on real IBM Cleveland hardware across diverse MedMNIST datasets, this work paves the way for future advancements in QML and highlights its potential to complement classical ML in the years to come.

### v2

@@ 2.4 Inference on real IBM Cleveland hardware
The trained quantum models demonstrating the best classification performance on the simulator were selected for validation on the IBM Cleveland quantum hardware. During the inference stage of our QML framework, the trained quantum models with optimized parameters were utilized for evaluating the test samples. For each testing sample, the corresponding quantum circuit is dynamically constructed and transpiled at optimization level 3 to minimize circuit depth and gate count while ensuring compatibility with the hardware topology. Moreover, qubit connectivity was carefully configured to align with the hardware’s topology, minimizing the need for SWAP gates, which can introduce unnecessary noise and overhead. To further enhance the quantum model performance, we employed error suppression techniques, i.e. dynamical decoupling(DD) and gate twirling (Twir) during the sampling process on real IBM hardware. The DD is used to suppress the decoherence effects introduced due to the unwanted interaction between the qubits and the environment. Gate twirling applies random single-qubit gates around a target gate to average out coherent errors. The sampling on real hardware is performed using Qiskit SamplerV2 in Session execution mode. Moreover, we also integrated M3 error mitigation [42] in our QML workflow. M3 mitigation is a scalable quantum error correction approach that operates within the noisy bitstring subspace, correcting correlated and uncorrelated errors using a memory-efficient, matrix-free iterative solver. The integration of these error suppression and mitigation strategies enables quantum models to achieve optimal inference performance on real quantum hardware. Lastly, classical post-processing is applied on the mitigated bitstrings to obtain the predicted classification labels. To achieve this, the mitigated bitstrings are processed to compute qubit probabilities defined as:
@@
where $\langle Z_{q}\rangle$ denotes the expectation value of the Pauli-Z operator for the $q$-th qubit, calculated as the difference between the probabilities of measuring 0 and 1. This expectation value vector obtained for each test sample is processed to obtain the predicted class labels. The results from all test samples were aggregated to compute the performance metrics, providing insights into the performance of the quantum circuits on real quantum hardware. By conducting inference on real quantum device, we demonstrated the feasibility of deploying noise-resilient and device-aware quantum circuits for practical medical image classification tasks.
@@ 3 Experimental results
3.1 Dataset and evaluation metrics details
In this subsection, we provide the details of the MedMNIST datasets [1] used in this benchmark study and performance metrics used to evaluate the QML workflow. The MedMNIST contains 12 two-dimensional biomedical imaging datasets of different medical modalities and complexity levels. These 12 datasets include 2-class as well as multi-class datasets including 4-class, 5-class, 7-class, 8-class, 9-class, and 11-class. For this study, we considered 8 datasets out of the 12 by making sure that at least one dataset from each class category was considered as shown in Table 1. The different medical modalities considered in these 8 datasets are illustrated in Fig. 2. We considered two widely used classification metrics i.e., accuracy (ACC) and area under the curve (AUC) to measure the performance of QML models. It is important to note that ACC provides a straightforward interpretation of model performance but it is highly sensitive to class imbalance, often leading to biased evaluations. On the contrary, AUC offers a more comprehensive evaluation of the classifier’s separability and is inherently robust to class imbalance. This makes AUC particularly suitable for datasets where class distributions are skewed. It is clear from Table 1 that the MedMNIST datasets are quite imbalanced. Therefore, we prioritized AUC over ACC to select a model for real hardware execution. By focusing on AUC, we ensured that the chosen quantum models were able to effectively distinguish between classes across various thresholds, providing a fairer and more reliable assessment of performance.
@@ 3.1 Dataset and evaluation metrics details
The classical post-processing of M3-mitigated bitstrings in the inference stage provides an expectation values vector for all test samples. These expectation values are used to calculate the ACC and AUC metrics. For binary datasets (where we measure one qubit), the accuracy is calculated by measuring the proportion of predictions that deviate from the true labels by less than a threshold of 1. It evaluates how many predictions are close enough to the actual labels, providing an overall measure of correctness for the model’s outputs in binary classification tasks. The binary AUC is computed by assessing the separability of the two classes on the basis of the predicted expectation values. We generate signed prediction labels to align with the true labels and calculate the AUC score using the AUC score function [43] from scikit-learn. For the multi-class datasets (where we are measuring more than 2 qubits), the model’s predicted class is determined by taking the index of the maximum value along each row of the expectation values vector. This is compared against the true labels to compute the classification accuracy as:
@@
To calculate the AUC metric for multi-class datasets, we first applied the softmax activation function to convert the raw expectation values into class probabilities. This ensures that the outputs for each test sample sum to 1 and can be interpreted as probabilities for each class. Then, we use the AUC function [43] from scikit-learn to calculate the AUC.
Table 2: Performance analysis of QML framework on simulator based on classification accuracy (ACC) and area under the curve (AUC) metrics under different configurations of quantum models.
@@
Table 3: Performance analysis of QML framework on simulator based on classification accuracy (ACC) and area under the curve (AUC) metrics under different configurations of quantum models.
@@ 3.2 Noiseless results
We present the noiseless results of the QML framework using a simulator based on ACC and AUC metrics under different settings of input embedding features and quantum circuit trainable parameters for the MedMNIST datasets we explored in Tables 2 and 3. The QML evaluation is performed by considering all the training and testing samples of the MedMNIST datasets as given in Table 1. The bold values in Tables 2 and 3 represent the best accuracy and AUC values. For the PneumoniaMNIST dataset, our QML model with 4 qubits, 49 input embedding features, and 120 trainable parameters provides the highest classification accuracy of 85.26% and AUC of 81.97% as compared to the other settings as shown in Table 2. In the case of the BreastMNIST dataset, we got the best accuracy of 81.40% for the 4-qubit quantum model having 49 embedding features and 100 trainable parameters, while the highest AUC of 72.20% is achieved for the model with 64 embedding features and 80 trainable parameters. Similarly, we got the highest accuracy of 44.30% and AUC of 65.90% for an 8-qubit quantum model having 49 embedding features and 100 trainable parameters for the OCTMNIST dataset. For the RetinaMNIST dataset, we achieved the best classification accuracy of 51.75% for the 5-qubit quantum model having 64 input embedding features and 120 trainable parameters. The highest AUC of 66.48% is obtained for the quantum model having 64 embedding features and 60 trainable parameters as depicted in Table 2. A 7-qubit quantum model with 64 input embedding features and 120 trainable parameters provides the highest accuracy of 67.38% and AUC of 73.43% for the 7-class DermaMNIST dataset as shown in Table 3. For an 8-class BloodMNIST dataset, we achieved the highest accuracy of 48.20% and AUC of 82.10% using the QML model with 64 embedding features and 120 trainable parameters. Similarly, we got the best accuracy of 37.23% and AUC of 78.40% for a 9-class PathMNIST dataset using the QML model having 64 embedding features and 120 trainable parameters. Lastly, the QML model with 64 features and 100 trainable parameters provides the best classification accuracy of 37.86% for the 11-class OrganSMNIST dataset, while the best AUC of 78.02% is achieved for the QML model having 64 embedding features and 120 trainable parameters. It is clear from these results that the quantum models without any classical neural networks and with very limited input features show promising capability to classify medical imaging datasets. Also, it is observed from the Tables 2 and 3 that there is a significant difference between the accuracy and AUC values, for instance, we got an accuracy of 48.20% and AUC of 82.10% for the BloodMNIST dataset. A similar trend can be seen for the PathMNIST and OrganSMNIST datasets. Note that we considered AUC as our primary performance evaluation metric because it is less sensitive to the class imbalance of the MedMNIST datasets as compared to the accuracy metric. Moreover, it can be noted that for most of the multi-class datasets, we got the best results corresponding to the QML models having more embedding features and trainable parameters in the quantum circuit.
3.3 Ablation study of error suppression and mitigation
We also performed an ablation study on IBM Cleveland to understand the impact of error suppression and error mitigation techniques on the performance of QML models for medical imaging classification as shown in Table 4. We performed this ablation on the OCTMNIST dataset because of the significant effect of hardware noise on the performance of the QML model. For the OCTMNIST dataset, our QML workflow provides the noiseless accuracy of 44.30% and AUC of 65.90% as shown in Table 2. When we ran this quantum model on real IBM hardware without any error suppression and mitigation techniques, we got an ACC of 28.60% and an AUC of 54.50%. Therefore, we observed a reduction of 15.70% and 11.40% in accuracy and AUC, respectively due to the noise of IBM Cleveland hardware. In this ablation, since we are using a sampler in the inference stage as shown in Fig. 1, we considered two error suppression techniques i.e., dynamical decoupling and gate twirling and one error mitigation technique i.e., M3 mitigation. It is observed from the Table 4 that when we use both error suppression techniques and M3 mitigation together (DD+Twirling+M3), we got the best accuracy of 39.70% and AUC of 62.20%, thereby providing an improvement of 11.10% and 7.7% in accuracy and AUC respectively when compared to the case where no suppression and mitigation is used. Based on this ablation study on the OCTMNIST dataset, we used this combination (DD+Twirling+M3) for all the MedMNIST datasets we studied.
@@ 3.3 Ablation study of error suppression and mitigation
3.4 IBM Cleveland hardware results
We selected the best performing QML models from Tables 2 and 3 based on the AUC metric for inference on IBM Cleveland for all considered MedMNIST datasets. Due to hardware limitations, we considered 1000 testing samples to evaluate the QML models for the datasets having a large number of testing samples. We presented noiseless and IBM hardware results (with and without error suppression and mitigation) together in Table 5 along with the specifications of the quantum models such as number of qubits, number of embedded features/gates, trainable parameters, quantum circuit gates, and circuit depth. It is observed from Table 5 that even with the small number of input embedding features, our QML models consistently provide promising results in terms of classification accuracy and AUC metrics. It is worth noting that the noiseless results and hardware results (without error suppression and mitigation) are quite comparable for most of the MedMNIST datasets except the OCTMNIST and OrganSMNIST datasets. This is because we generated device-aware quantum circuits with optimal circuit gates, depth, and robustness against hardware noise. For the OCTMNIST dataset, the classification accuracy and AUC on the hardware (without error suppression and mitigation) are reduced by 15.70% and 11.40% respectively as compared to the noiseless results. Similarly, accuracy and AUC decreased by 14.11% and 4.8% respectively for the OrganSMNIST dataset as compared to the noiseless results. This indicates that there is scope for further improvement in the design of device-aware quantum circuits. To mitigate the effect of hardware noise, we ran the hardware experiments for the selected MedMNIST datasets using an error suppression and error mitigation combination (DD+Twirling+M3) based on the ablation study conducted in Table 4. We noted improvements in classification accuracy and AUC results using the DD+Twirling+M3 combination as compared to the hardware results without error suppression and mitigation for all the datasets except the BreastMNIST dataset. For instance, the classification accuracy and AUC values are significantly improved by 11.10% and 7.7% respectively as compared to the hardware results without error suppression and mitigation for the OCTMNIST dataset. For the BreastMNIST dataset, the accuracy and AUC values decreased by 1.3% and 2.4% respectively when compared to the hardware results without error suppression and mitigation.
@@ 3.4 IBM Cleveland hardware results
Table 6: Comparative analysis of QML model and classical benchmark ML models using AUC and ACC metrices on MedMNIST datasets.
PneumoniaMNIST BreastMNIST OCTMNIST RetinaMNIST
Method (Input size) AUC ACC AUC ACC AUC ACC AUC ACC
ResNet-18 (28$\times$28$\times$3) [44] 0.944 0.854 0.901 0.863 0.943 0.743 0.717 0.524
ResNet-18 (224$\times$224$\times$3) [44] 0.956 0.864 0.891 0.833 0.958 0.763 0.71 0.493
ResNet-50 (28$\times$28$\times$3) [44] 0.948 0.854 0.857 0.812 0.952 0.762 0.726 0.528
ResNet-50 (224$\times$224$\times$3) [44] 0.962 0.884 0.866 0.842 0.958 0.776 0.716 0.511
Auto-sklearn (28$\times$28$\times$3) [45] 0.942 0.855 0.836 0.803 0.887 0.601 0.69 0.515
AutoKeras (28$\times$28$\times$3) [46] 0.947 0.878 0.871 0.831 0.955 0.763 0.719 0.503
Google AutoML
Vision (28$\times$28$\times$3) [47]
0.991 0.946 0.919 0.861 0.963 0.771 0.75 0.531
QML model (7$\times$7 or 8$\times$8)
(IBM Cleveland)
0.822 0.854 0.714 0.769 0.622 0.397 0.664 0.505
DermaMNIST BloodMNIST PathMNIST OrganSMNIST
Method (Input size) AUC ACC AUC ACC AUC ACC AUC ACC
ResNet-18 (28$\times$28$\times$3) [44] 0.917 0.735 0.998 0.958 0.983 0.907 0.972 0.782
ResNet-18 (224$\times$224$\times$3) [44] 0.92 0.754 0.998 0.963 0.989 0.909 0.974 0.778
ResNet-50 (28$\times$28$\times$3) [44] 0.913 0.735 0.997 0.956 0.99 0.911 0.972 0.77
ResNet-50 (224$\times$224$\times$3) [44] 0.912 0.731 0.997 0.95 0.989 0.892 0.975 0.785
Auto-sklearn (28$\times$28$\times$3) [45] 0.902 0.719 0.984 0.878 0.934 0.716 0.945 0.672
AutoKeras (28$\times$28$\times$3) [46] 0.915 0.749 0.998 0.961 0.959 0.834 0.974 0.813
Google AutoML
Vision (28$\times$28$\times$3) [47]
0.914 0.768 0.998 0.966 0.944 0.728 0.964 0.749
QML model (7$\times$7 or 8$\times$8)
(IBM Cleveland)
0.698 0.677 0.826 0.424 0.78 0.418 0.768 0.34
3.5 Comparison with classical ML models
@@
We also provided a comparative analysis with the benchmarks of the classical ML models provided in MedMNIST [1] as shown in Table 6. To the best of our knowledge, this benchmarking study is the first to provide a comprehensive evaluation of QML (without using any classical neural networks) on real IBM quantum hardware using large-scale MedMNIST datasets. Although the classification results obtained by our QML workflow do not outperform the best-performing classical ML model due to the inherent limitations of current NISQ hardware, our results represent a significant step forward in evaluating QML for realistic applications in medical imaging. All classical ML models leverage the full spatial dimensions of input images, whereas our QML model operates on a significantly reduced feature set, dictated by quantum hardware constraints. Despite this, our QML model achieves promising results on IBM Cleveland in terms of classification accuracy and AUC values. For the 2-class PneumoniaMNIST dataset, we obtained good AUC and classification accuracies of 82.2% and 85.4% respectively. Our QML accuracy of 85.4% for the PneumoniaMNIST dataset is equal to the accuracy obtained by the classical ML models including ResNet-18 (28) [44] and ResNet-50 (28) [44] as shown in Table 6. Also, our QML accuracy 85.4% for PneumoniaMNIST dataset is lagging by only 0.1%, 1%, 2.4%, and 3% when compared with the Auto-sklearn [45], ResNet-18 (224) [44], AutoKeras [46], and ResNet-50 (224) [44] classical ML models, respectively. Similarly, our QML model classification accuracy is lagging by 3.4% when compared to the Auto-sklearn [45] classical ML model for the BreastMNIST dataset. Moreover, in the case of the 5-class RetinaMNIST dataset, we achieved an AUC and accuracy of 66.4% and 50.5% respectively which lags by 8.6% and 2.6% in AUC and accuracy when compared to the best performing classical Google AutoML Vision [47] model. The AUC value 66.4% of our QML model for RetinaMNIST lags only by 2.6% compared to the Auto-sklearn [45] classical ML model. Importantly, our QML model accuracy of 50.5% for RetinaMNIST outperforms the ResNet-18 (224) [44] and AutoKeras [46] classical ML models by 1.2% and 0.2%, respectively. Even for the 7-class DermaMNIST dataset, our QML model obtained good results with AUC and accuracy values of 69.8% and 67.7% respectively. The accuracy value achieved by our QML model for the DermaMNIST dataset only lags by 4.2% and 9.1% when compared to the Auto-sklearn [45] model and the best performing classical Google AutoML Vision [47] ML model, respectively. Our QML model can achieve an AUC of more than 70% for all the MedMNIST datasets we explored except for the OCTMNIST, RetinaMNIST, and DermaMNIST datasets. It is worth mentioning that for the 2-class PneumoniaMNIST and 8-class BloodMNIST datasets, we obtained AUCs of more than 80%. These results underscore the promise of QML in medical image classification, particularly when benchmarked against the existing literature on image classification using quantum approaches, where achieving comparable performance with classical models remains an open challenge. Our study provides a critical evaluation of QML for realistic applications, showcasing its capability even in the current phase of quantum computing. By systematically benchmarking QML performance on IBM Cleveland across diverse MedMNIST datasets, this work paves the way for future advancements in QML and highlights its potential to complement classical ML in the years to come.
To further assess the effectiveness of our QML model, we conducted a comparative analysis by training classical ML baselines, ResNet-18 and ResNet-50, using the same 8×8 input image features as used in our QML model. This ensures a fair and direct comparison of the 5-class RetinaMNIST dataset, as summarized in Table 7. We selected the RetinaMNIST dataset for this experiment because it presents a high level of complexity, resulting in the lowest performance among all classical ML models. Despite having significantly fewer trainable parameters (60) compared to ResNet-18 (11.2 million) and ResNet-50 (23.5 million), our QML model demonstrates superior classification performance. Our QML model surpasses ResNet-18 and ResNet-50 by improving classification accuracy by 8% and 7%, respectively, while also achieving AUC gains of 5.41% and 1.54%, respectively. These results highlight that our QML model achieves quantum advantage in both performance and computational complexity. Motivated by these findings, we aim to further explore advanced embedding techniques to incorporate more features into quantum models, as well as to further optimize ansatz design to enhance expressivity. These improvements will be crucial in demonstrating quantum advantage on a larger scale in future experiments.
Table 7: Comparative analysis of QML model and classical benchmark ML models by considering same number of input features on 5-class RetinaMNIST dataset.
RetinaMNIST (5-class)
Method (Input size) Trainable parameters ACC AUC
ResNet-18 (8$\times$8) [44] 11.2 million 0.425 0.6101
ResNet-50 (8$\times$8) [44] 23.5 million 0.435 0.6488
QML model (8$\times$8)
(IBM Cleveland)
60 0.505 0.6642